Light travelling in air strikes a flat piece of uniformly thick glass at an incident angle of 60.0°, as shown. If the index of refraction of the glass is 1.50, (a) what is the angle of refraction θA in the glass; (b) what is the angle θB at which the ray emerges from the glass?


1
Expert's answer
2021-07-07T08:39:21-0400


According to Snell's Law

"n_1sin \theta_1= n_2sin \theta_2"

a) When ray strikes the thick mirror

"n_1=1 \\

n_2=1.5 \\

\theta_1=60^o \\

\theta_2=\theta_A \\

1 \times sin60 =1.5 \times sin \theta_A \\

\theta_A =35.26^o"

b) when the light emerges out from thick mirror

"n_1=1.5 \\

n_2=1 \\

\theta_1=35.26^o \\

\theta_2=\theta_B \\

1.5 \times sin 35.26=1 \times sin \theta_B \\

\theta_B= 60^o"